### v2

# Many-body quantum geometry in time-dependent quantum systems with emergent quantum field theory instantaneously

Abstract
We study many-body quantum geometric effects in time-dependent system with emergent quantum integrable field theory instantaneously. We establish a theorem stating that the Berry connection matrix thus all associated geometric quantities of the system can be precisely characterized by excitations up to two particles from the initial quantum integrable system. To illustrate the many-body geometric influence, we analyze an Ising chain subjected to both a small longitudinal field and a slowly rotating transverse field, whose low-energy physics in the scaling limit is instantaneously governed by the quantum $E_{8}$ integrable field theory. Focusing on the quantum geometric potential (QGP), we show the QGP continuously suppresses the instantaneous energy gaps with decreasing longitudinal field, thereby enhancing many-body Landau-Zener tunneling as evidenced by the Loschmidt echo and its associated spectral entropy. The critical threshold for the longitudinal field strength is determined, where the spectral entropy linearly increases with system size and exhibits hyperscaling behavior when approaching to the threshold. As the longitudinal field passes the threshold and decreases toward zero, the QGP continuously leads to vanishing instantaneous energy gaps involving more low-energy excitations, resulting in increasing spectral entropy indicative of many-body Landau-Zener tunneling. Our results unveil telltale quantum geometric signatures in time-dependent many-body systems, elucidating the intricate interplay between quantum geometry and dynamics.
Introduction.— Quantum geometric effects have attracted sustained research interests over the past several decades since the introduction of the Berry phase [1]. Significant progress has been made in fields such as topological band insulators [2, 3, 4], non-equilibrium adiabatic evolution [5, 6], and the quantum Hall effect [7]. For instance, the higher-order optical responses can be generally characterized by the Riemannian geometry of quantum states [2, 3, 4]. Despite the clear structure of quantum geometric quantities constructed from wavefunctions, studying quantum geometric effects in quantum many-body systems remains highly challenging, and related research is still scarce [8]. This scarcity primarily stems from the complexity of many-body wavefunctions which involve an exponentially growing Hilbert space in such systems. For instance, calculating Berry connections in many-body systems poses significant challenges, both numerically and analytically.
Due to exact solvability, quantum integrable systems may provide critical insights into quantum geometry in many-body systems. For most quantum integrable systems, the eigenstates and corresponding transition matrix elements of physical observables can be calculated exactly using systematic methods, such as the Bethe Ansatz  [9, 10, 11, 12, 13, 14, 15, 16, 17, 18] or/and form factor bootstrap [19, 20, 21, 22, 23]. Over the past decade, significant research has focused on quantum integrability in non-equilibrium systems, including integrable quantum quenches and integrable brickwork circuits [24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36]. These studies have provided new insights into non-equilibrium physics from the perspective of integrability, with potential connections to emerging fields such as quantum computing and quantum information. However, the understanding of quantum geometric effects in quantum many-body systems remains underdeveloped [37].
To study quantum geometry and its effects in quantum many-body systems, we focus on time-dependent systems with emergent quantum integrable field theory instantaneously. The $n^{th}$ instantaneous eigenfunction is $|\phi_{n}(t)\rangle$ with instantaneous eigenenergy $e_{n}(t)$ at the moment $t$. After imposing a proper constraint on the adiabatic transformation $\mathscr{U}_{a}(t):|{\phi}_{n}(0)\rangle\to|{\phi}_{n}(t)\rangle$, we prove that the Berry connection matrix (BCM) entries $\gamma_{nm}=i\langle\phi_{n}(t)|\dot{\phi}_{m}(t)\rangle$ only associate excitations up to two particle processes.
For illustration we further consider the time evolution of a quantum Ising chain under both a slowly rotating transverse field and a small static longitudinal field [38, 35], where the initial state is set as the ground state at the beginning. We introduce a local U(1) gauge-invariant geometric term known as the quantum geometric potential (QGP) $Q_{mn}=\gamma_{mm}-\gamma_{nn}+\frac{d}{dt}\text{arg}~{}\gamma_{nm}$  [39, 40] generated from any two instantaneous eigenfunctions. This term provides a crucial correction to the instantaneous energy gap $e_{nm}=e_{n}(t)-e_{m}(t)$ and thus gives rise to a more efficient effective energy gap (EEG) $\Delta_{nm}=e_{nm}+Q_{mn}$. Notably, even when the instantaneous energy gap $e_{nm}$ is large, perfect Landau-Zener tunneling can still occur if the $\Delta_{nm}$ vanishes when the QGP cancels the $e_{nm}$. This has been validated in the NMR experiment [41] and further demonstrated in time-dependent two-band models [5].
In the scaling limit, the QGP of this system can be analytically determined through the quantum $E_{8}$ integrable field theory (IFT). The EEG can then be reduced to a more compact formalism as a function of the rotating frequency ($\omega_{0}$) and the longitudinal field intensity ($h_{z}$). We first identify our calculation of the QGP with the truncated conformal space approach (TCSA).
Then to show the influence of the QGP and EEG in many-body Landau-Zener tunneling (MBLZT) processes, we employ the truncated lattice free fermion approach (TLFFA) [42, 43, 44, 45]. Within this framework, we analyze the spectral entropy of the Loschmidt echo (LE), which quantifies the overlap between the time-evolved wavefunction and the instantaneous ground state for different system sizes $L$. We find a critical $\kappa\equiv h_{z}/(\omega_{0}/2)$, $\kappa_{c}=1$ which serves as a critical threshold for the MBLZT. When $\kappa>1$, zeros of EEG are nearly absent, manifested as normal spectral entropy contributed from a limited number of finite-frequency modes. As $\kappa$ goes to 1, a hyperscaling behavior of spectral entropy appears according to the scaling parameter $(\kappa-1)^{8/15}L$. At $\kappa=1$, the spectral entropy scales linearly with the system size $L$, implying MBLZT avalanche in thermodynamic limit. When $\kappa\leq 1$ zeros of EEG substantially increase with decreasing $\kappa$, manifested as persistent growth of the LE spectral entropy, signaling significantly enhanced MBLZT. Interestingly, the LE spectral entropy also exhibits a cascade of additional peaks when $\kappa\leq 1$, where we expect strengthened MBLZT.
Figure 1: An illustration of an eigenstate in an IFT containing $n$ stable quasiparticles on a cylinder of a circumference $L$, with the adiabatic transformation $\mathscr{U}_{a}(t)$.
Figure 2: (a) Comparison of the QGP between the ground state (vacuum of the $\mathcal{H}_{E_{8}}$) and the three lightest single-particle states of the $\mathcal{H}_{E_{8}}$, calculated with both the $\mathcal{H}_{E_{8}}$ form factor theory and the boosted TCSA algorithm  [46]. Here, $\Lambda$ is the truncation parameter satisfying $E^{2}-P^{2}\leq\Lambda^{2}$. (b) Analytical results of the EEG between the ground state and the 8 single $E_{8}$ particle state with zero momentum, with $0<\kappa\leq 2$. The zeros of EEG appear at $\kappa=8/15$ for all the single $E_{8}$ particles. (c) The zeros of EEG as a function of $A_{1}$’s rapidity $\vartheta_{1}$, for the two-particle states $A_{1}A_{1}$ to $A_{1}A_{4}$, with a zero momentum constraint $m_{i}\sinh\vartheta_{1}+m_{j}\sinh\vartheta_{2}=0$. (d) Zeros of EEG between two four $E_{8}$ particle states. While both the two groups of particles are of particle type $A_{1\sim 4}$, the rapidities for one of the states takes $(\vartheta_{1},\vartheta_{2},\vartheta_{3},\vartheta_{4})=(0.2,0.2,0.1,0.1)$, and for the other one takes $\vartheta_{1}=\vartheta_{2}=\vartheta_{4}=0.1$ and $\vartheta_{3}=0.2421,0.2422,0.2425$, corresponding to the zeros being from large to small, respectively.
The BCM in instantaneously quantum-integrable time-dependent systems.— It is generally difficult to calculate the BCM in interacting quantum many-body systems. We consider time-dependent quantum system with emergent 1+1D quantum field theory instantaneously, where the initial state is an eigenstate in an IFT containing $n$ stable quasiparticles, and the instantaneous eigenstate is connected to the initial state with an adiabatic transformation, as $|\phi_{n}(t)\rangle=\mathscr{U}_{a}(t)|\phi_{n}(0)\rangle_{\text{init},L}$. Without loss of generality the system is placed on a cylinder with circumference $L$. We require the effective adiabatic Hamiltonian [47] $\mathcal{H}_{adia}=i\mathscr{U}_{a}^{-1}(t)\dot{\mathscr{U}_{a}}(t)$ to satisfy
[…]
in the continuum limit. Here $\mathcal{O}_{i}(x)$ are the local field densities. The initial state is given by $|\phi_{n}(0)\rangle_{\text{init},L}=|A_{1}(\vartheta_{1};I_{1})...A_{N}(\vartheta_{N};I_{N})\rangle_{L}$. Here, $A_{j},j=1,...,N$ label different types of particles with corresponding mass $m_{j}$, discrete quantum number $I_{j}$, and rapidity $\vartheta_{j}$. In addition, the energy and momentum of this state are given by $E=\sum_{j}m_{j}\cosh\vartheta_{j}$ and $P=\sum_{j}m_{j}\sinh\vartheta_{j}$. For any single particle $A_{j}(\vartheta_{j};I_{j})$, the discrete momentum and rapidity are connected with each other by the Bethe-Yang quantization condition derived from the spatial periodic boundary condition [48, 49, 50, 35, 42]. An illustration is shown in Fig. 1. The BCM can then be explicitly given with the help of quantum integrability. Considering two such states containing $n$ and $m$ quasiparticles respectively, for the off-diagonal matrix element $m\neq n$, the BCM is given by
$\frac{\gamma_{nm}}{L}\approx\sum_{i}f_{i}(t)\frac{\mathcal{F}_{NM}^{\mathcal{O}_{i}}(\vartheta^{\prime}_{1}+i\pi,...,\vartheta^{\prime}_{N}+i\pi,\vartheta_{1},...,\vartheta_{M})}{\sqrt{\rho_{N}(\vartheta^{\prime}_{1},...,\vartheta^{\prime}_{N})}\sqrt{\rho_{M}(\vartheta_{1},...,\vartheta_{M})}},$ (2)
with Lüscher’s remainder $o(e^{-\mu L})$ [51, 48, 49, 50]. Here, the function $\mathcal{F}_{nm}^{\mathcal{O}_{i}}$ in terms of rapidities is known as the form factor for certain IFT, and the $\rho$ function comes from the Jacobian of the particle density of states determined by the Bethe-Yang condition. For the diagonal matrix element $m=n$, the result is given by
$\frac{\gamma_{nn}}{L}$ $\approx\sum_{i}f_{i}(t)\bigg{[}{}_{L}\langle\mathcal{O}_{i}\rangle_{L}+$ (3)
$\sum_{A\in\{A_{1},...,A_{N}\}}\frac{\mathcal{F}_{NN}^{\mathcal{O}_{i}}(A)\rho(\{A_{1},...,A_{N}\}\setminus A)_{L}}{\rho(\{A_{1},...,A_{N}\})_{L}}\bigg{]},$
with the Lüscher’s remainder as well. The full information of the BCM implies that for such quantum many-body systems, we can determine all the geometric quantities associated with the BCM. Notably, the Jacobians in the denominator provide suppression behavior of order $L^{n}$ with $n$ being the number of particles involved. Consequently, after simple power counting, we can immediately observe that the excitations involving more than two particles do not contribute to the BCM in the thermodynamic limit ($L\rightarrow\infty$). Thus we prove the following theorem.
Theorem: Consider time-dependent many-body quantum systems with emergent quantum integrable field theory instantaneously, if the effective adiabatic Hamiltonian follows Eq. (1), then in the thermodynamic limit the corresponding BCM entries can be analytically obtained as excitations up to two particles.
In the following, the theorem is applied to a time-dependent quantum Ising model, where the QGP constructed from the BCM entries is shown to play a crucial role in the MBLZT in the system.
A time-dependent quantum Ising chain with instantaneous quantum $E_{8}$ integrability.— We start from a ferromagnetic Ising chain subjected to both a periodically rotating transverse field and a static longitudinal field  [38, 35],
$\mathcal{H}(t)=-\sum_{i=1}^{L}\left(\sigma^{z}_{i}\sigma^{z}_{i+1}+\cos\omega_{0}t\sigma^{x}_{i}-\sin\omega_{0}t\sigma^{y}_{i}+h_{z}\sigma^{z}_{i}\right),$ (4)
with $\hbar=c=1$, Pauli matrices $\sigma_{i}^{\alpha}(\alpha=x,~{}y,~{}z)$ at site $i$, the length of the chain $L$, and the neighboring spin coupling $J=1$. Additionally, we require $\omega_{0}$, $h_{z}\ll 1$. It is not hard to observe that the Hamiltonian instantaneously appears as a quantum critical transverse-field Ising chain perturbed by a small longitudinal field, whose low-energy physics in the scaling limit is described by the quantum $E_{8}$ IFT,
[…]
where $\mathcal{H}_{c=1/2}$ is the central charge $1/2$ conformal field theory, $\sigma(x)$ is known as the spin density operator, and the coupling constant $h$ gives an exact relation with the mass of the lightest particle as $m_{1}=C_{m}h^{8/15}$, with $C_{m}=4.4049$…, and masses of all quantum $E_{8}$ particles in units of $m_{1}$ satisfy the eigenvalues of the Cartan matrix for $E_{8}$ Lie algebra [52, 22, 35, 53, 54]. In the following we set the initial state $|\psi(0)\rangle$ as the ground state of Eq. (4) at $t=0$, which in the scaling limit is the vacuum state $|0\rangle_{h}$ of the $\mathcal{H}_{E_{8}}$.
The general time-evolved wavefunction of the Hamiltonian Eq. (4) follows [38, 35] $|\psi(t)\rangle=\mathcal{U}(t)^{-1}e^{-i\mathcal{H}_{\text{eff}}t}|\psi(0)\rangle$, where the unitary transformation is given by $\mathcal{U}(t)=\exp\left\{-\frac{i}{2}\omega_{0}t\sum_{i}\sigma_{i}^{z}\right\}$, satisfying $\mathcal{U}(t)\mathcal{H}(t)\mathcal{U}(t)^{-1}=\mathcal{H}(t=0)$, and the effective Hamiltonian $\mathcal{H}_{\text{eff}}=\mathcal{U}\mathcal{H}(t)\mathcal{U}^{-1}+i\dot{\mathcal{U}}\mathcal{U}^{-1}$ follows,
$\mathcal{H}_{\text{eff}}=-\sum_{i}\left[\sigma_{i}^{z}\sigma_{i+1}^{z}+\sigma_{i}^{x}+(h_{z}-\frac{\omega_{0}}{2})\sigma_{i}^{z}\right],$ (6)
where for $h_{z}\neq{\omega_{0}}/{2}$, the low-energy physics is described by another $\mathcal{H}_{E_{8}}$ with modified longitudinal coupling. Considering the instantaneous Schrödinger equation $\mathcal{H}(t)|\psi_{n}(t)\rangle_{ins}=E_{n}(t)|\psi_{n}(t)\rangle_{ins}$, where $|\psi_{n}(t)\rangle_{ins}$ is the instantaneous eigenstate associated with the $n^{th}$ instantaneous eigenenergy $E_{n}(t)$, it is straightforward to prove [42] $E_{n}(t)=E_{n}(0)$ and $|\psi_{n}(t)\rangle_{ins}=e^{i\varphi_{n}(t)}\mathcal{U}(t)^{-1}|\psi_{n}(0)\rangle$, which implies that the instantaneous energy levels remain invariant for the Hamiltonian Eq. (4), and the instantaneous eigenstates are given by a global rotation according to $\mathcal{U}(t)$ up to an additional time-dependent overall U(1) gauge $\varphi_{n}(t)$.
Now let’s focus on the QGP $Q_{mn}$, taking the scaling limit of the model Eq. (4), the QGP follows
$Q_{nm}=\sum_{j=1}^{M}\frac{\mathcal{F}^{\mathcal{O}}_{jj}(i\pi,0)}{m_{j}\cosh\vartheta_{j}}-\sum_{i=1}^{N}\frac{\mathcal{F}^{\mathcal{O}}_{ii}(i\pi,0)}{m_{i}\cosh\vartheta_{i}}.$ (7)
As a concrete example, we consider the QGP between the vacuum state and the single $E_{8}$ particle state with zero momentum, which is given by $\mathcal{A}_{00}-\mathcal{A}_{kk}=\langle\sigma\rangle\mathcal{F}^{\sigma}_{kk}(i\pi,0)/m_{k}$ where $\langle\sigma\rangle=C_{\sigma}h^{1/15}$ with $C_{\sigma}=1.27758\cdots$ [22, 35]. The results agree with those from the boosted TCSA algorithm [46], as shown in Fig. 2 (a) for $k=1,2,3$.
The EEG in time-dependent IFT.— In the following we consider the EEG between two instantaneous eigenstates and try to find the zeros in $\Delta_{mn}(t)=0$ for any $|\psi_{m}\rangle_{ins}$ and $|\psi_{n}\rangle_{ins}$. At the zeros, the quantum adiabatic condition is significantly violated [39, 41, 40, 5], and the MBLZT process shall occur. Moreover, when $h_{z}$ decreases the EEG will cross zero and become negative, which suggests an effective swap of the high- and low- energy states instantaneously. Under this scenario, the MBLZT will be strongly enhanced because more and more higher-energy states effectively become low-energy due to the QGP. In the following we analytically solve the EEG for the ”low-energy” sector of the Hamiltonian Eq. (4), whose results are associated with the $\mathcal{H}_{E_{8}}$ [Eq. (5)].
[…]
$Q_{nm}(t)=-\frac{8}{15\kappa}\left(\sum_{\beta=1}^{M}\frac{m_{\beta}}{\cosh\theta_{\beta}}-\sum_{\alpha=1}^{N}\frac{m_{\alpha}}{\cosh\theta^{\prime}_{\alpha}}\right),$ (8)
with $\kappa\equiv h_{z}/(\omega_{0}/2)$ as introduced in the introduction. Thus the EEG finally reads
$\Delta_{mn}(t)=$ $\sum_{\beta=1}^{M}m_{\beta}\cosh\theta_{\beta}-\sum_{\alpha=1}^{N}m_{\alpha}\cosh\theta^{\prime}_{\alpha}$ (9)
$-\frac{8}{15\kappa}\left(\sum_{\beta=1}^{M}\frac{m_{\beta}}{\cosh\theta_{\beta}}-\sum_{\alpha=1}^{N}\frac{m_{\alpha}}{\cosh\theta^{\prime}_{\alpha}}\right).$
Interestingly, if taking one of the states being the ground state (vacuum) of the initial states, and the other one being the single particle state with zero momentum, the zeros of EEG are given by $\kappa=8/15$, which does not depend on the type of the particles [cf. Fig. 2 (b)]. It is also the upper bound for the zeros of EEG between the ground state and the higher excited states [42]. In the quantum $E_{8}$ case, the higher excited states are readily given by single particle states with large rapidity (momentum) or multi-particle states. We show that for two-particle excitations with zero total momentum constraint, as one of the particles’ rapidity goes to infinity, the corresponding zeros of the EEG will approach 0. It is also easy to verify from Eq. (9) that the zeros for two same particle combinations with zero total momentum will collapse to the same curve as $A_{1}A_{1}$. The results are presented in Fig. 2 (c). However, if considering the zeros of EEG for any two instantaneous eigenstates, the region for zeros of EEG is beyond $\kappa=8/15$. For instance, Fig. 2 (d) shows contribution of four-particle excitations to the EEG whose zeros appear to approach $\kappa=1$.
The spectral entropy of the Loschmidt echo.— In order to investigate the role of the EEG in the MBLZT process of a time-dependent quantum system, a commonly adopted approach is to examine the fidelity between the time-evolving wavefunction and the instantaneous eigenstates, denoted as ${}_{ins}\langle\psi_{n}(t)|\psi(t)\rangle$, which is known as the Loschmidt echo (LE).
For the time evolution governed by Eq. (4), the LE is expressed as $\mathcal{L}_{n}(t)=\langle\psi_{n}(0)|e^{-i\mathcal{H}_{\text{eff}}t}|\psi(0)\rangle$. Here, the index $n$ is used to label the energy level of the corresponding excited states. When $n$ corresponds to the ground state, we focus on the spectrum of the LE obtained through its Fourier transformation, which is given by $\mathcal{L}_{0}(\omega)=\sum_{m}|\langle\phi_{m}|\psi_{0}\rangle|^{2}\delta(\omega-E_{m})$. Here, $|\phi_{m}\rangle$ represents the eigenbasis of the $\mathcal{H}_{\text{eff}}$, and $E_{m}$ is its corresponding eigenenergy. Detailed results for $L=80$ with different $h_{z}$ from $\omega_{0}/10$ to $\omega_{0}$ are exhibited in [42]. To obtain a better understanding of the MBLZT, we introduce spectral entropy for the LE,
$S_{\text{LE}}=-\sum_{m}p_{m}\ln p_{m},$ (10)
where $p_{m}=|\langle\phi_{m}|\psi_{0}\rangle|^{2}$, and it naturally satisfies $\sum_{m}p_{m}=1$. In the context of our study, the $S_{\text{LE}}$ serves as a measure of the spectral density across the entire spectrum.
[…]
The entire process can also be qualitatively understood in terms of the $\mathcal{H}_{\text{eff}}$ [Eq. (6)]. Precisely at $\kappa=1$, the $\mathcal{H}_{\text{eff}}$ represents a quantum critical transverse field Ising chain with divergent density of states in the low energy sector, leading to a significant increase of the spectral entropy. When $\kappa<1$, the spin alignment in the ground state of $\mathcal{H}_{\text{eff}}$ is opposite to that of the initial ground state. This means that the initial state becomes a high-energy state with respect to the $\mathcal{H}_{\text{eff}}$, resulting in progressively stronger MBLZT as $\kappa$ decreases.
Conclusions.— In this article we lay down a theorem revealing the quantum geometry in time-dependent many-body systems whose low-energy can be organized by quantum integrable field theory instantaneously. The theorem states that the BCM entries for local field densities in the system are governed by at most two-particle excitations. The result provides a geometric foundation for diabatic time evolutions mediated by the vanishing EEG due to the QGP contribution. We then analyze a time-dependent quantum Ising chain, showing that the suppression of the EEG with decreasing longitudinal field enhances the MBLZT. Near the critical ratio $\kappa=1$, the hyperscaling behavior of the LE spectral entropy signals a sharp transition in the MBLZT. For $\kappa<1$, the numerical results show persistent entropy growth mixing with a cascade of peaks, reflecting continuous enhancement of MBLZT. Our results indicate that the LE spectral entropy can serve as a powerful diagnostic tool for diabatic time evolution and its associated MBLZT. Yet, a full physical understanding on the analytical structure of the LE spectral entropy, such as its hyperscaling behavior, linear in $L$ behavior at $\kappa=1$, and additional peaks when $\kappa<1$, requires more analytical and numerical efforts, a challenge that remains open but worth to explore further. Meanwhile it is also worth to generalize the theorem in quantum integrable lattice systems, and extend to the BCM entries for quasi-local field densities.
The time-dependent Hamiltonian in Eq. (4) is experimentally accessible in platforms such as Rydberg atomic arrays, [55, 56, 35] a promising setup for direct experimental verification of our theoretical predictions. Furthermore, our study reveals the important role of quantum geometry in time-dependent quantum many-body systems, providing an important insight in understanding non-equilibrium many-body systems.
Figure 3: $S_{LE}$ $vs.$ $\kappa$ for different system sizes $L$’s. The $S_{LE}$ at $\kappa=1$ linearly increases with $L$ [inset (a)], along with more peaks emerging for $\kappa\lesssim 8/15$. When $\kappa$ approaches 1, the $S_{LE}$ collapses on the same curve for different $L$ [inset (b)], implying hyperscaling behavior of $S_{\text{LE}}$ w.r.t scaling parameter $(\kappa-1)^{8/15}L$.
Acknowledgements.— We thank Zongping Gong, Qicheng Tang, Thors Hans Hansson and Gabor Takács for helpful discussions. The TCSA calculation is supported by the TCSA package in [46]. The work is sponsored by National Natural Science Foundation of China Nos. 12274288, 12450004, the Innovation Program for Quantum Science and Technology Grant No. 2021ZD0301900, and the Tsung-Dao Lee Scholarship.
Appendix A Supplemental Material—Many-body quantum geometry in time-dependent quantum systems with emergent quantum field theory instantaneously
Appendix B A. Finite Size formalism for Integrable Field Theory
In this section, we briefly introduce the framwork of integrable field theory (IFT) that is defined on a cylinder with a sufficiently large spatial length $L$, while time ranges from $\infty$- ∞ to +$\infty$+ ∞. The formalism has been well developed by B. Pozsgay and G. Takács, and more detailed reviews can be found in [48, 49, 50]. The asymptotic free multi-particle state for an infinite IFT is denoted by the particle species and their corresponding rapidities, as
$|\psi_{N}\rangle=|A_{a_{1}}(\vartheta_{1})...A_{a_{N}}(\vartheta_{N})\rangle,$ (11)
with the normalization as $\langle A_{a_{i}}(\theta_{i})|A_{a_{j}}(\theta_{j})\rangle=2\pi\delta_{a_{i}a_{j}}\delta(\theta_{i}-\theta_{j})$. In the finite size framework, such multi-particle states are relabeled with the quantum numbers as
$|\psi_{N}\rangle_{L}=|A_{a_{1}}(I_{1})...A_{a_{N}}(I_{N})\rangle_{L},$ (12)
with the normalization as $\langle A_{a_{i}}(I_{i})|A_{a_{j}}(I_{j})\rangle=\delta_{a_{i}a_{j}}\delta_{I_{i}I_{j}}$. Due to the periodic boundary conditions, the quantum numbers are connected with the rapidities via the Bethe-Yang equation from the finite size field theory scheme,
$m_{i}L\sinh\theta_{i}-i\sum_{j\neq i}\ln S_{ij}(\theta_{i}-\theta_{j})=2\pi I_{i}=Q_{i},$ (13)
with $S_{ij}(\theta_{i}-\theta_{j})$ being the corresponding scattering matrix. However, it is clearly observed that there is a measure difference between the normalizations of the two quantum states. Thus all the related quantities such as the complete basis and form factors (FF) will be influenced. This issue can be done by introducing the Jacobian given by the corresponding particle sets,
$\rho_{a_{1},...,a_{n}}(\theta_{1},...,\theta_{n})=\text{Det}\left(\frac{\partial Q_{k}(\theta_{1},...,\theta_{n})}{\partial\theta_{l}}\right),$ (14)
and we have the following measure transformation between all the associated quantities,
0$\langle$ 0⟨ 0 $|\mathcal{O}|A_{a_{1}}(I_{1})...A_{a_{N}}(I_{N})\rangle_{L}=\frac{F^{\mathcal{O}}_{n}(\theta_{1},...,\theta_{n})_{a_{1},...,a_{n}}}{\sqrt{\rho_{a_{1},...,a_{n}}(\theta_{1},...,\theta_{n})}},$ (15)
$|A_{a_{1}}(I_{1})...A_{a_{N}}(I_{N})\rangle_{L}=\frac{|A_{a_{1}}(\theta_{1})...A_{a_{N}}(\theta_{N})\rangle}{\sqrt{\rho_{a_{1},...,a_{n}}(\theta_{1},...,\theta_{n})}},$
$\sum_{I_{1},...,I_{n}}=\int\frac{d\theta_{1}...d\theta_{n}}{(2\pi)^{n}}\rho_{a_{1},...,a_{n}}(\theta_{1},...,\theta_{n}),$
where $F^{\mathcal{O}}_{n}(\theta_{1},...,\theta_{n})_{a_{1},...,a_{n}}=\langle 0|\mathcal{O}|A_{a_{1}}(\theta_{1})...A_{a_{N}}(\theta_{N})\rangle$ is known as the (elementary) FF for operator $\mathcal{O}$.
Appendix C B. A proof to the theorem
This section gives the proof of the theorem in the main text. Considering an adiabatic transformation of an eigenstate in an IFT containing $n$ stable quasiparticles,
$|\psi_{n}(t)\rangle=\mathscr{U}(t)|\psi_{n}\rangle_{\text{init},L},$ (16)
where we have put this quantum field theory on a cylinder with circumference $L$, with the adiabatic transformation satisfying
[…]
which is also known as the effective adiabatic Hamiltonian [47]. The initial state is given by
$|\psi_{n}\rangle_{\text{init},L}=|A_{1}(\vartheta_{1};I_{1})...A_{N}(\vartheta_{N};I_{N})\rangle_{L}.$ (18)
The related parameters have been introduced in the main text. According to the definition of the Berry connections, we focus on calculating the quantities $\gamma_{nm}=i\langle\psi_{n}(t)|\dot{\psi}_{m}(t)\rangle$. Introducing $|\psi_{m}(t)\rangle$ with $M$ stable quasi-particles, then we have [50]
$\gamma_{nm}=L\sum_{i=1}^{N_{\mathcal{O}}}f_{i}(t)\langle\psi_{n}|\mathcal{O}_{i}|\psi_{m}\rangle|_{p_{\psi_{n}}=p_{\psi_{m}}},$ (19)
where $p_{\psi_{n}}$ and $p_{\psi_{m}}$ are the total momentum of the corresponding quantum states. Using the framework listed in Sec. A, let’s consider a specific $\mathcal{O}_{i}$, using the crossing relation [22, 50, 53, 35], we can transfer the matrix element for a general FF to the elementary FF as listed in Eq. (15),
$F_{nm}^{\mathcal{O}}(\theta^{\prime}_{1},...,\theta^{\prime}_{n}|\theta_{1},...,\theta_{m})$ $=\langle A_{1}(\vartheta^{\prime}_{1})...A_{n}(\vartheta^{\prime}_{n})|\mathcal{O}|A_{1}(\vartheta_{1})...A_{m}(\vartheta_{m})\rangle$ (20)
$=F_{n-1,m+1}^{\mathcal{O}}(\theta^{\prime}_{1},...,\theta^{\prime}_{n-1}|\theta^{\prime}_{n}+i\pi,\theta_{1},...,\theta_{m})+\sum_{k=1}^{m}\Bigg{(}2\pi\delta_{a_{n}b_{k}}\delta(\theta^{\prime}_{n}-\theta_{k})\prod_{l=1}^{k-1}S_{b_{l}b_{k}}(\theta_{l}-\theta_{k})$
×$\times F^{\mathcal{O}}_{n-1,m-1}(\theta^{\prime}_{1},...,\theta^{\prime}_{n-1}|\theta_{1},...,\theta_{k-1},\theta_{k+1},...,\theta_{m})\Bigg{)}.$
Then for the off-diagonal matrix element $n\neq m$, we have
$\langle\psi_{n}|\mathcal{O}_{i}|\psi_{m}\rangle={}_{L}\langle A_{1}(\vartheta^{\prime}_{1};I^{\prime}_{1})...A_{N}(\vartheta^{\prime}_{N};I^{\prime}_{N})|\mathcal{O}_{i}|A_{1}(\vartheta_{1};I_{1})...A_{M}(\vartheta_{M};I_{M})\rangle_{L}=\frac{\mathcal{F}_{NM}^{\mathcal{O}_{i}}(\vartheta^{\prime}_{1}+i\pi,...,\vartheta^{\prime}_{N}+i\pi,\vartheta_{1},...,\vartheta_{M})}{\sqrt{\rho_{N}(\vartheta^{\prime}_{1},...,\vartheta^{\prime}_{N})}\sqrt{\rho_{M}(\vartheta_{1},...,\vartheta_{M})}},$ (21)
More interestingly, the diagonal matrix element is known to be the disconnected terms and show a divergence due to the delta function. A way to correctly recover such divergence in the finite size framework of IFT is developed in [48, 49, 50], which follows,
${}_{L}\langle A_{1}(\vartheta_{1};I_{1})...A_{M}(\vartheta_{M};I_{M})|\mathcal{O}_{i}|A_{1}(\vartheta_{1};I_{1})...A_{M}(\vartheta_{M};I_{M})\rangle_{L}\approx{}_{L}\langle\mathcal{O}_{i}\rangle_{L}+\sum_{A\in\{A_{1},...,A_{N}\}}\frac{\mathcal{F}_{NN}^{\mathcal{O}_{i}}(A)\rho(\{A_{1},...,A_{N}\}\setminus A)_{L}}{\rho(\{A_{1},...,A_{N}\})_{L}},$ (22)
where $A$ is particle set contains particles with types $a_{1}$ to $a_{M}$, $\mathcal{F}$ is FF corresponding to particle set $A$, By making use of Eq. (21) and Eq. (22), we obtain the results for the Berry connection matrix entries. Finally, for the off-diagonal matrix entries $m\neq n$,
$\frac{\gamma_{nm}}{L}\approx\sum_{i}f_{i}(t)\frac{\mathcal{F}_{NM}^{\mathcal{O}_{i}}(\vartheta^{\prime}_{1}+i\pi,...,\vartheta^{\prime}_{N}+i\pi,\vartheta_{1},...,\vartheta_{M})}{\sqrt{\rho_{N}(\vartheta^{\prime}_{1},...,\vartheta^{\prime}_{N})}\sqrt{\rho_{M}(\vartheta_{1},...,\vartheta_{M})}}.$ (23)
For the diagonal matrix elements $m=n$,
$\frac{\gamma_{nn}}{L}\approx\sum_{i}f_{i}(t)\bigg{[}{}_{L}\langle\mathcal{O}_{i}\rangle_{L}+\sum_{A\in\{A_{1},...,A_{N}\}}\frac{\mathcal{F}_{NN}^{\mathcal{O}_{i}}(A)\rho(\{A_{1},...,A_{N}\}\setminus A)_{L}}{\rho(\{A_{1},...,A_{N}\})_{L}}\bigg{]},$ (24)
as shown in the main text. Noticing that the Jacobian Eq. (14) is a polynomial of order $L^{M}$, it is easy to see that in Eq. (23) only the term that $(M+N)/2\leq 1$ [as in Eq. (24) only the term that $N\leq 1$] finally survives in the thermodynamic limit, suggesting that the contributions for the Berry connection matrix elements are up to two particle excitations.
Appendix D C. The instantaneous energy levels and eigenstates
We investigate the instantaneous Schrödinger equation $\mathcal{H}(t)|\psi_{n}(t)\rangle_{ins}=E_{n}(t)|\psi_{n}(t)\rangle_{ins}$ for the following Hamiltonian,
$\mathcal{H}(t)=-J\sum_{i=1}^{L}\left(\sigma^{z}_{i}\sigma^{z}_{i+1}+\cos\omega_{0}t\sigma^{x}_{i}-\sin\omega_{0}t\sigma^{y}_{i}+h_{z}\sigma^{z}_{i}\right).$ (25)
Under the unitary transformation $\mathcal{U}(t)=\exp\left\{-\frac{i}{2}\omega_{0}t\sum_{i}\sigma_{i}^{z}\right\}$, we have [38, 35]
$\mathcal{U}(t)\mathcal{H}(t)\mathcal{U}(t)^{-1}=-J\sum_{i=1}^{L}\left(\sigma^{z}_{i}\sigma^{z}_{i+1}+\sigma^{x}_{i}+h_{z}\sigma^{z}_{i}\right)=\mathcal{H}_{t=0}.$ (26)
Applying the eigenstate of $\mathcal{H}_{t=0}$ on both side,
$\mathcal{U}(t)\mathcal{H}(t)\mathcal{U}(t)^{-1}|\psi_{n}(0)\rangle=\mathcal{H}_{t=0}|\psi_{n}(0)\rangle=E_{n}|\psi_{n}(0)\rangle,$ (27)
and this further gives
$\mathcal{H}(t)\mathcal{U}(t)^{-1}|\psi_{n}(0)\rangle=E_{n}\mathcal{U}(t)^{-1}|\psi_{n}(0)\rangle,$ (28)
i.e., the instantaneous eigenstate $|\psi_{n}(t)\rangle_{ins}$ is proportional to $\mathcal{U}(t)^{-1}|\psi_{n}(0)\rangle$ while they have the same time-independent eigenenergy $E_{n}$. Without loss of generality, we conclude
$E_{n}(t)=E_{n}(0),~{}~{}|\psi_{n}(t)\rangle_{ins}=e^{i\varphi_{n}(t)}\mathcal{U}(t)^{-1}|\psi_{n}(0)\rangle,$ (29)
with $\varphi_{n}(t)$ being a time-dependent U(1) gauge.
Appendix E D. Calculation of the quantum geometric potential
The effective energy gap is defined as $\Delta_{mn}(t)=E_{m}(t)-E_{n}(t)+Q_{nm}(t)$, where $E_{m,n}(t)$ are the instantaneous eigenenergy of the corresponding states, and
$Q_{nm}(t)=\gamma_{nn}(t)-\gamma_{mm}(t)+\frac{d}{dt}\text{arg}\left[\langle\psi_{m}(t)|\dot{\psi_{n}}(t)\rangle\right]_{ins}$ (30)
is known as the quantum geometric potential (QGP), with $\gamma_{nn}(t)={}_{ins}\langle\psi_{n}(t)|i\partial_{t}|\psi_{n}(t)\rangle_{ins}$ being the Berry connection [39, 40, 5]. We consider the effective energy gap between any two arbitrary excited states, which in the scaling limit corresponds to two multi-$E_{8}$ particle states containing $N$ and $M$ quantum $E_{8}$ particles, respectively. The instantaneous energy gap is proven to be static, and can be directly read as
$E_{m}(t)-E_{n}(t)=\sum_{\beta=1}^{M}m_{\beta}\cosh\theta_{\beta}-\sum_{\alpha=1}^{N}m_{\alpha}\cosh\theta^{\prime}_{\alpha}.$ (31)
On the lattice, using Eq. (29), the Berry connection is given by
$\gamma_{mm}(t)$ $=-\dot{\varphi}_{m}(t)-\frac{\omega_{0}}{2}\sum_{i=1}^{L}\langle\psi_{m}(0)|\sigma^{z}_{i}|\psi_{m}(0)\rangle,$ (32)
$\gamma_{nn}(t)$ $=-\dot{\varphi}_{n}(t)-\frac{\omega_{0}}{2}\sum_{i=1}^{L}\langle\psi_{n}(0)|\sigma^{z}_{i}|\psi_{n}(0)\rangle.$
Since we also have
$\frac{d}{dt}\text{arg}\left[\langle\psi_{m}(t)|\dot{\psi_{n}}(t)\rangle\right]_{ins}=\dot{\varphi}_{n}(t)-\dot{\varphi}_{m}(t).$ (33)
The QGP is given as following,
$Q_{nm}(t)=\frac{\omega_{0}}{2}\sum_{i=1}^{L}\left(\langle\psi_{m}(0)|\sigma^{z}_{i}|\psi_{m}(0)\rangle-\langle\psi_{n}(0)|\sigma^{z}_{i}|\psi_{n}(0)\rangle\right),$ (34)
which also implies that the QGP is U(1) gauge invariant. To obtain the exact expression of the QGP, we use the finite size field theory frame and the results in Sec. A and B [48, 49, 50]. We first transform the system into the scaling limit. Since $\langle\sigma\rangle=\bar{s}J^{1/8}\langle\sigma^{z}\rangle$ [57, 35],
$Q^{s}_{nm}=\frac{\omega}{2}\bar{s}^{-1}J^{-1/8}\lim_{L\rightarrow\infty}2JL\left({}_{L}\langle\theta_{1}...\theta_{M}|\sigma|\theta_{1}...\theta_{M}\rangle_{L}-{}_{L}\langle\theta^{\prime}_{1}...\theta^{\prime}_{N}|\sigma|\theta^{\prime}_{1}...\theta^{\prime}_{N}\rangle_{L}\right),$ (35)
with $\omega$ being the scaling limit of $\omega_{0}$, $L=Na$ being the length of the system, and $a$ being the lattice spacing, satisfying $12Ja=\hbar c=$ 1 [22, 53]. Introducing the general FF for the quantum $E_{8}$ field theory,
$F^{\mathcal{O}}_{a_{1},...,a_{n},b_{1},...,b_{m}}(\theta_{1},...,\theta_{n},\theta_{n+1},...,\theta_{n+m})={}_{a_{1},...,a_{n}}\langle\theta_{1},...,\theta_{n}|\mathcal{O}|\theta_{n+1},...,\theta_{n+m}\rangle_{b_{1},...,b_{m}}/\langle\mathcal{O}\rangle.$ (36)
The $E_{8}$ FF theory and a complete FF bootstrapping process can be found in Refs. [22, 50, 53]. According to previous analysis of the FFs, the QGP follows,
$Q^{s}_{nm}(t)=\frac{\omega}{2}\bar{s}^{-1}J^{-1/8}\langle\sigma\rangle 2J\left(\sum_{\beta=1}^{M}\frac{F^{\sigma}_{\beta\beta}(i\pi,0)}{m_{\beta}\cosh\theta_{\beta}}-\sum_{\alpha=1}^{N}\frac{F^{\sigma}_{\alpha\alpha}(i\pi,0)}{m_{\alpha}\cosh\theta^{\prime}_{\alpha}}\right).$ (37)
This result implies that the QGP for this system is only related with the diagonal two-particle FFs in quantum $E_{8}$ field theory. In fact, it can be further reduced to a more transparent form. Since $m_{\beta}=C^{s}_{m}h^{8/15}e_{\beta}$, with $e_{\beta}$ being the Cartan roots of $E_{8}$ exceptional Lie algebra, $\langle\sigma\rangle=C^{s}_{\sigma}h^{1/15}$, with $h$ being $h_{z}$ in the scaling limit, and $C^{s}_{\sigma}=4(C^{s}_{m})^{2}/15\varphi_{11}$, where $\varphi_{11}$ is a constant [22], we can obtain that,
$Q^{s}_{nm}(t)$ $=\frac{\omega}{2}\bar{s}^{-1}J^{-1/8}\langle\sigma\rangle 2J\left(\sum_{\beta=1}^{M}\frac{F^{\sigma}_{\beta\beta}(i\pi,0)}{m_{\beta}\cosh\theta_{\beta}}-\sum_{\alpha=1}^{N}\frac{F^{\sigma}_{\alpha\alpha}(i\pi,0)}{m_{\alpha}\cosh\theta^{\prime}_{\alpha}}\right)$ (38)
$=\frac{\omega}{2}\frac{4C^{s}_{m}}{15\varphi_{11}}\bar{s}^{-1}J^{-1/8}2Jh^{-7/15}\left(\sum_{\beta=1}^{M}\frac{F^{\sigma}_{\beta\beta}(i\pi,0)}{m_{\beta}\cosh\theta_{\beta}}-\sum_{\alpha=1}^{N}\frac{F^{\sigma}_{\alpha\alpha}(i\pi,0)}{m_{\alpha}\cosh\theta^{\prime}_{\alpha}}\right).$
It can be proven that $F^{\sigma}_{\beta\beta}(i\pi,0)/e^{2}_{\beta}=-2\varphi_{11}$, which is not related with specific particle type. For showing this, we consider the trace of the stress-energy tensor $\Theta(x)$ [22],
$\langle m_{\beta}|\Theta(x)|m_{\beta}\rangle=2\pi m_{\beta}^{2}=2\pi(C^{s}_{m})^{2}h^{16/15}e_{\beta}^{2}.$ (39)
On the other hand, since $\Theta(x)=2\pi h(2-2\Delta_{\sigma})\sigma(x)$, where $\Delta_{\sigma}=1/16$ is the conformal weight of $\sigma(x)$, together with $C^{s}_{\sigma}=4(C^{s}_{m})^{2}/15\varphi_{11}$, we can immediately obtain that $F^{\sigma}_{\beta\beta}(i\pi,0)/e^{2}_{\beta}=-2\varphi_{11}$. Then the QGP is further reduced as
$Q^{s}_{nm}(t)$ $=-\frac{\omega}{2}\frac{8}{15h}\bar{s}^{-1}J^{-1/8}2J\left(\sum_{\beta=1}^{M}\frac{m_{\beta}}{\cosh\theta_{\beta}}-\sum_{\alpha=1}^{N}\frac{m_{\alpha}}{\cosh\theta^{\prime}_{\alpha}}\right).$ (40)
Sending back to the lattice limit, using $h=2J^{15/8}h_{z}/\bar{s}$ [53, 35], we finally obtain that
$Q_{nm}(t)=-\frac{\omega_{0}}{2J}\frac{8}{15h_{z}}\left(\sum_{\beta=1}^{M}\frac{m_{\beta}}{\cosh\theta_{\beta}}-\sum_{\alpha=1}^{N}\frac{m_{\alpha}}{\cosh\theta^{\prime}_{\alpha}}\right).$ (41)
This implies that the QGPs in this system are only related with the contributions from all the quantum $E_{8}$ particles in the corresponding quantum $E_{8}$ states.
Appendix F E. Technical Details of TLFFA
We provide additional details about the truncated lattice free fermion approach (TLFFA). Given the Hamiltonian
$\mathcal{H}=\mathcal{H}_{0}-Jh_{z}\sum_{i=1}^{L}\sigma_{i}^{z},$ (42)
where $\mathcal{H}_{0}$ represents the Hamiltonian of the transverse field Ising chain with periodic boundary condition,
$\mathcal{H}_{0}=-J\sum_{i=1}^{L}\left(\sigma^{z}_{i}\sigma^{z}_{i+1}+g\sigma^{x}_{i}\right).$ (43)
To construct the Hamiltonian for TLFFA, we introduce the dual Jordan-Wigner (JW) transformation  [43]. The usual JW transformation is given by [58, 59]
$\sigma_{i}^{z}=\prod_{i^{\prime}<i}\left(1-2c_{i^{\prime}}^{\dagger}c_{i^{\prime}}\right)\left(c_{i}^{\dagger}+c_{i}\right),~{}~{}\sigma_{i}^{x}=1-2c_{i}^{\dagger}c_{i}.$ (44)
Thus the Hamiltonian will be transformed into
$\mathcal{H}_{0}=-J\sum_{i=1}^{L}\left[\left(c_{i}^{\dagger}-c_{i}\right)\left(c_{i+1}^{\dagger}+c_{i+1}\right)-g\left(c_{i}^{\dagger}-c_{i}\right)\left(c_{i}^{\dagger}+c_{i}\right)\right].$ (45)
The dual JW transformation is then given by [43]
$c_{i}^{\dagger}-c_{i}=-(a_{i}^{\dagger}-a_{i}),~{}~{}c_{i+1}^{\dagger}+c_{i+1}=a_{i}^{\dagger}+a_{i},$ (46)
with $a_{j}^{\dagger}$ and $a_{j}$ being dual JW fermion creation and annihilation operators, and $\mathcal{H}_{0}$ follows
$\mathcal{H}_{0}=2J\sum_{i=1}^{L}\left(a_{i}^{\dagger}a_{i}-\frac{1}{2}\right)-Jg\sum_{i=1}^{L}\left(a_{i+1}^{\dagger}-a_{i+1}\right)\left(a_{i}^{\dagger}+a_{i}\right).$ (47)
The corresponding Hilbert space is expressed as a direct sum of the Ramond (R, $\mathcal{V}_{R}$) and Neveu-Schwarz (NS, $\mathcal{V}_{NS}$) sectors. We introduce the Bogoliubov transformation [59]
$\gamma_{q}=u_{q}a_{q}-iv_{q}a_{-q}^{\dagger},~{}~{}\gamma_{q}^{\dagger}=u_{q}a_{q}^{\dagger}+iv_{q}a_{-q},$ (48)
with
$u_{q}=\cos\frac{\theta_{q}}{2},~{}~{}v_{q}=\sin\frac{\theta_{q}}{2},~{}~{}\tan\theta_{q}=\frac{g\sin q}{1-g\cos q},$ (49)
in particular for $q=0$,
$u_{q}=1,~{}~{}v_{q}=0.$ (50)
$\mathcal{H}_{0}$ can then be diagonalised as
$\mathcal{H}_{0}=\sum_{q}\varepsilon_{q}\left(\gamma_{q}^{\dagger}\gamma_{q}-\frac{1}{2}\right),$ (51)
where
$\varepsilon_{q}=2J\sqrt{1+g^{2}-2g\cos q}.$ (52)
The eigenvectors in the two subspaces are
$|P\rangle_{R}=|p_{1},\dots,p_{n}\rangle_{R}=\gamma_{p_{1}}^{\dagger}\gamma_{p_{2}}^{\dagger}\gamma_{p_{3}}^{\dagger}\dots\gamma_{p_{n}}^{\dagger}|0\rangle_{R}\in\mathcal{V}_{R},$ (53)
$|Q\rangle_{NS}=|q_{1},\dots,q_{m}\rangle_{NS}=\gamma_{q_{1}}^{\dagger}\gamma_{q_{2}}^{\dagger}\gamma_{q_{3}}^{\dagger}\dots\gamma_{q_{m}}^{\dagger}|0\rangle_{NS}\in\mathcal{V}_{NS},$ (54)
with the quantum number $p$ and $q$
$\mathcal{P}:\left\{p=\frac{2\pi}{L}j\right\},~{}~{}\mathcal{Q}:\left\{q=\frac{2\pi}{L}\left(j-\frac{1}{2}\right)\right\},~{}~{}j=-\frac{L}{2}+1,-\frac{L}{2}+2,\dots,\frac{L}{2}.$ (55)
The non-trivial part for the construction of Eq. (42) is given by the FF of $\sigma^{z}$ operator, and the FF of $\sigma^{z}_{1}$ follows [43]
${}_{R}{\langle P|}\sigma_{1}^{z}|Q\rangle_{NS}$ $={}_{R}{\langle p_{1},p_{2},\dots,p_{n}|}\sigma_{1}^{z}|q_{1},q_{2},\dots,q_{m}\rangle_{NS}$ (56)
$=\delta_{m-n,0}^{(\text{mod}\ 2)}i^{-(n+m)/2}(-1)^{n(n-1)/2}\prod_{p\in P}e^{-ip/2}\prod_{q\in Q}e^{iq/2}\left(\frac{2}{L}\right)^{(m+n)/2}h^{(m-n)^{2}/4}\sqrt{\xi\xi_{T}}$
×$\quad\times\prod_{q\in Q}\frac{e^{\eta(q)/2}}{\sqrt{\varepsilon_{q}}}\prod_{p\in P}\frac{e^{-\eta(p)/2}}{\sqrt{\varepsilon_{p}}}\prod_{q<q^{\prime}\in Q}\frac{4\sin\left(\frac{q-q^{\prime}}{2}\right)}{\varepsilon_{q}+\varepsilon_{q^{\prime}}}\prod_{p<p^{\prime}\in P}\frac{4\sin\left(\frac{p-p^{\prime}}{2}\right)}{\varepsilon_{p}+\varepsilon_{p^{\prime}}}\prod_{q\in Q}\prod_{p\in P}\frac{\varepsilon_{q}+\varepsilon_{p}}{4\sin\left(\frac{q-p}{2}\right)},$
where $\xi=(1-g^{2})^{1/4}$ and
$\xi_{T}=\frac{\prod_{q\in\mathcal{Q}}\prod_{p\in\mathcal{P}}(\varepsilon_{q}+\varepsilon_{p})^{1/2}}{\prod_{q,q^{\prime}\in\mathcal{Q}}(\varepsilon_{q}+\varepsilon_{q^{\prime}})^{1/4}\prod_{p,p^{\prime}\in\mathcal{P}}(\varepsilon_{p}+\varepsilon_{p^{\prime}})^{1/4}},~{}~{}e^{\eta(\alpha)}=\frac{\prod_{q^{\prime}\in\mathcal{Q}}(\varepsilon_{\alpha}+\varepsilon_{q^{\prime}})}{\prod_{p^{\prime}\in\mathcal{P}}(\varepsilon_{\alpha}+\varepsilon_{p^{\prime}})}.$ (57)
The spin operator $\sigma_{l}^{z}$ is then associated with $\sigma_{1}^{z}$ as
${}_{R}{\langle P|}\sigma_{l}^{z}|Q\rangle_{NS}=e^{i(l-1)\left(\sum_{q\in Q}q-\sum_{p\in P}p\right)}{}_{R}{\langle P|}\sigma_{1}^{z}|Q\rangle_{NS}.$ (58)
We can then obtain all the matrix elements for Eq. (42). However, due to the exponential growth of the Hilbert space, a truncation of the full Hamiltonian is necessary, which is the key to TLFFA. In this paper, we include only states that satisfy the zero total momentum constraint and have an energy upper bound of $10J$ [44, 45],
$\sum_{i}k_{i}=0\;{\text{and}}\;\sum_{k_{i}}\varepsilon_{k_{i}}\leq 10J.$ (59)
Figure 4: $-p\ln p$ $vs.$ $\omega$ for different $\kappa$ from $0.2$ to $2$ with $\Delta\kappa=0.2$, corresponding to the curves from the bottom to the top, after a Lorentzian broadening with $\alpha=0.001$.
Appendix G F. LE spectrum
The detailed data for the LE spectral entropy $S_{\text{LE}}=-\sum_{n}p_{n}\ln p_{n}$, where $p_{n}=|\langle\phi_{n}|\psi_{0}\rangle|^{2}$ are provided. These data were calculated using the TLFFA algorithm with $L=80$ and 8192 truncated states for several values of $\kappa$. It can be observed that as $\kappa$ decreases to $\kappa<1$, the number of peaks increases, accompanied by spectral weight growth. This suggests an increase in the LE spectral entropy and, consequently, an enhancement of the many-body Landau-Zener tunneling due to the quantum geometric effects.