### v3

# Berry connection and quantum geometry in time-dependent systems with instantaneous quantum integrable field theory

Abstract
We study many-body quantum geometric effects in time-dependent system with emergent quantum integrable field theory instantaneously. We establish a theorem stating that the Berry connection matrix thus all associated geometric and topological quantities of the system can be precisely characterized by excitations up to two particles from the initial emergent quantum integrable field theory. To illustrate the many-body geometric influence, we analyze an Ising chain subjected to both a small longitudinal field and a slowly rotating transverse field, whose low-energy physics in the scaling limit is instantaneously governed by the quantum $E_{8}$ integrable field theory. Focusing on the quantum geometric potential (QGP), we show the QGP continuously suppresses the instantaneous energy gaps with decreasing longitudinal field, thereby enhancing many-body Landau-Zener tunneling as evidenced by the Loschmidt echo and its associated spectral entropy. The critical threshold for the longitudinal field strength is determined, where the spectral entropy linearly increases with system size and exhibits hyperscaling behavior when approaching to the threshold. As the longitudinal field passes the threshold and decreases toward zero, the QGP continuously leads to vanishing instantaneous energy gaps involving more low-energy excitations, resulting in increasing spectral entropy indicative of many-body Landau-Zener tunneling. Our results unveil telltale quantum geometric signatures in time-dependent many-body systems, elucidating the intricate interplay between quantum geometry, integrability, and dynamics.
Introduction.— Quantum geometric effects have attracted sustained research interests over the past several decades since the introduction of the Berry phase [1]. Significant progress has been made in fields such as topological band insulators [2, 3, 4], non-equilibrium adiabatic evolution [5, 6], and the quantum Hall effect [7]. For instance, the higher-order optical responses can be generally characterized by the Riemannian geometry of quantum states [2, 3, 4]. Despite the clear structure of quantum geometric quantities constructed from wavefunctions, studying quantum geometric effects in quantum many-body systems remains highly challenging, and related research is still scarce [8]. This scarcity primarily stems from the complexity of many-body wavefunctions which involve an exponentially growing Hilbert space in such systems. For instance, calculating Berry connections in many-body systems poses significant challenges, both numerically and analytically.
Due to exact solvability, quantum integrable systems may provide critical insights into quantum geometry in many-body systems: Their eigenstates can be completely labeled by the corresponding conserved quantum numbers [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20], therefore provide a concrete base to analytically study quantum geometry in quantum integrable systems [21, 22, 23, 24, 25, 26]. Over the past decade, significant research has focused on quantum integrability in non-equilibrium systems, including integrable quantum quenches and integrable brickwork circuits [27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39]. These studies have provided new insights into non-equilibrium physics from the perspective of integrability, with potential connections to emerging fields such as quantum computing and quantum information. However, the understanding of quantum geometric effects in quantum many-body (integrable) systems remains underdeveloped [40].
To study quantum geometry and its effects in quantum many-body systems, we focus on time-dependent systems with emergent quantum integrable field theory instantaneously. The $n$-th instantaneous eigenfunction is $|\phi_{n}(t)\rangle$ with instantaneous eigenenergy $e_{n}(t)$ at the moment $t$. After imposing a proper constraint on the adiabatic transformation $\mathscr{U}_{a}(t):|{\phi}_{n}(0)\rangle\to|{\phi}_{n}(t)\rangle$, we prove that the Berry connection matrix (BCM) entries $\gamma_{nm}=i\langle\phi_{n}(t)|\dot{\phi}_{m}(t)\rangle$ only associate excitations up to two particle processes.
For illustration we further consider the time evolution of a quantum Ising chain under both a slowly rotating transverse field and a small static longitudinal field [41, 38], where the initial state is set as the ground state at the beginning. We introduce a local U(1) gauge-invariant geometric term known as the quantum geometric potential (QGP) $Q_{mn}=\gamma_{mm}-\gamma_{nn}+\frac{d}{dt}\text{arg}~\gamma_{nm}$  [25, 26] generated from any two instantaneous eigenfunctions. This term provides a crucial correction to the instantaneous energy gap $e_{nm}=e_{n}(t)-e_{m}(t)$ and thus gives rise to a more efficient effective energy gap (EEG) $\Delta_{nm}=e_{nm}+Q_{mn}$. Notably, even when the instantaneous energy gap $e_{nm}$ is large, perfect Landau-Zener tunneling can still occur if the $\Delta_{nm}$ vanishes when the QGP cancels the $e_{nm}$. This has been validated in the NMR experiment [42] and further demonstrated in time-dependent two-band models [5].
In the scaling limit, the QGP of this system can be analytically determined through the quantum $E_{8}$ integrable field theory (IFT). The EEG can then be reduced to a more compact formalism as a function of the rotating frequency ($\omega_{0}$) and the longitudinal field intensity ($h_{z}$). We first identify our calculation of the QGP with the truncated conformal space approach (TCSA).
Then to show the influence of the QGP and EEG in many-body Landau-Zener tunneling (MBLZT) processes, we employ the truncated lattice free fermion approach (TLFFA) [43, 44, 45, 46]. Within this framework, we analyze the spectral entropy of the Loschmidt echo (LE), which quantifies the overlap between the time-evolved wavefunction and the instantaneous ground state for different system sizes $L$. We find a critical $\kappa\equiv h_{z}/(\omega_{0}/2)$, $\kappa_{c}=1$ which serves as a critical threshold for the MBLZT. When $\kappa>1$, zeros of EEG are nearly absent, manifested as normal spectral entropy contributed from a limited number of finite-frequency modes. As $\kappa$ goes to 1, a hyperscaling behavior of spectral entropy appears according to the scaling parameter $(\kappa-1)^{8/15}L$. At $\kappa=1$, the spectral entropy scales linearly with the system size $L$, implying MBLZT avalanche in thermodynamic limit. When $\kappa\leq 1$ zeros of EEG substantially increase with decreasing $\kappa$, manifested as persistent growth of the LE spectral entropy, signaling significantly enhanced MBLZT. Interestingly, the LE spectral entropy also exhibits a cascade of additional peaks when $\kappa\leq 1$, where we expect strengthened MBLZT.
Figure 1: An illustration of the setup for the theorem. (a) The initial wavefunction of an eigenstate in an IFT contains NN quasiparticles with periodic boundary conditions, represented as a cylinder with circumference $L$. The blue circles represent the quasiparticles. (b) The time evolution of the initial wavefunction under the unitary transformation $\mathscr{U}_{a}(t)$, with the effective adiabatic Hamiltonian given by Eq. (1). The blue line corresponds to the initial configuration in (a), and the red space-time network illustrates $\mathscr{U}_{a}(t)$.
Figure 2: (a) Comparison of the QGP between the ground state (vacuum of the $\mathcal{H}_{E_{8}}$) and the three lightest single-particle states of the $\mathcal{H}_{E_{8}}$, calculated with both the $\mathcal{H}_{E_{8}}$ form factor theory and the boosted TCSA algorithm  [47]. Here, $\Lambda$ is the truncation parameter satisfying $E^{2}-P^{2}\leq\Lambda^{2}$. (b) Analytical results of the EEG between the ground state and the 8 single $E_{8}$ particle state with zero momentum, with $0<\kappa\leq 2$. The zeros of EEG appear at $\kappa=8/15$ for all the single $E_{8}$ particles. (c) The zeros of EEG as a function of $A_{1}$’s rapidity $\vartheta_{1}$, for the two-particle states $A_{1}A_{1}$ to $A_{1}A_{4}$, with a zero momentum constraint $m_{i}\sinh\vartheta_{1}+m_{j}\sinh\vartheta_{2}=0$. (d) Zeros of EEG between two four $E_{8}$ particle states. While both the two groups of particles are of particle type $A_{1\sim 4}$, the rapidities for one of the states takes $(\vartheta_{1},\vartheta_{2},\vartheta_{3},\vartheta_{4})=(0.2,0.2,0.1,0.1)$, and for the other one takes $\vartheta_{1}=\vartheta_{2}=\vartheta_{4}=0.1$ and $\vartheta_{3}=0.2421,0.2422,0.2425$, corresponding to the zeros being from large to small, respectively.
The BCM in instantaneously quantum-integrable time-dependent systems.— It is generally difficult to calculate the BCM in interacting quantum many-body systems. We consider time-dependent quantum system with emergent 1+1D quantum field theory instantaneously, where the initial state is an eigenstate in an IFT containing $N$ stable quasiparticles [Fig. 1(a)], and the instantaneous eigenstate is connected to the initial state with an adiabatic transformation, as $|\phi_{N}(t)\rangle=\mathscr{U}_{a}(t)|\phi_{N}(0)\rangle_{\text{init},L}$ [Fig. 1(b)]. Without loss of generality the system is placed on a cylinder with circumference $L$. We require the effective adiabatic Hamiltonian [48] $\mathcal{H}_{adia}=i\mathscr{U}_{a}^{-1}(t)\dot{\mathscr{U}_{a}}(t)$ to satisfy
$\mathcal{H}_{adia}=\sum_{i=1}^{N_{\mathcal{O}}}\int_{0}^{L}f_{i}(t)\mathcal{O}_{i}(x)dx$ (1)
in the continuum limit. Here $\mathcal{O}_{i}(x)$ are the local field densities. The initial state is given by $|\phi_{N}(0)\rangle_{\text{init},L}=|A_{1}(\vartheta_{1};I_{1})...A_{N}(\vartheta_{N};I_{N})\rangle_{L}$. Here, $A_{j},j=1,...,N$ label different types of particles with corresponding mass $m_{j}$, discrete quantum number $I_{j}$, and rapidity $\vartheta_{j}$. In addition, the energy and momentum of this state are given by $E=\sum_{j}m_{j}\cosh\vartheta_{j}$ and $P=\sum_{j}m_{j}\sinh\vartheta_{j}$. For any single particle $A_{j}(\vartheta_{j};I_{j})$, the discrete momentum and rapidity are connected with each other by the Bethe-Yang quantization condition derived from the spatial periodic boundary condition [49, 50, 51, 38, 43]. The BCM can then be explicitly given with the help of quantum integrability. Considering two such states containing $N$ and $M$ quasiparticles respectively, for the off-diagonal matrix element where $M\neq N$, or $M=N$ but the two sets of rapidities are not exactly same, the BCM is given by
$\frac{\gamma_{NM}}{L}\approx\sum_{i}f_{i}(t)\frac{\mathcal{F}_{NM}^{\mathcal{O}_{i}}(\{\vartheta^{\prime}+i\pi\}_{N},\{\vartheta\}_{M})}{\sqrt{\rho_{L,\{\mathscr{A^{\prime}}\}_{N}}}\sqrt{\rho_{L,\{\mathscr{A}\}_{M}}}},$ (2)
with Lüscher’s remainder $o(e^{-\mu L})$ [52, 49, 50, 51]. Here, $\{\vartheta^{\prime}+i\pi\}_{N}=\{\vartheta^{\prime}_{1}+i\pi,...,\vartheta^{\prime}_{N}+i\pi$} and $\{\vartheta\}_{M}=\{\vartheta_{1},...,\vartheta_{M}\}$ are the sets of rapidities of the two particle sets $\{\mathscr{A^{\prime}}\}_{N}=\{A^{\prime}_{1},...,A^{\prime}_{N}\}$ and $\{\mathscr{A}\}_{M}=\{A_{1},...,A_{M}\}$. The function $\mathcal{F}_{NM}^{\mathcal{O}_{i}}(\{\vartheta^{\prime}+i\pi\}_{N},\{\vartheta\}_{M})$ in terms of rapidities is the matrix element of operator $\mathcal{O}_{i}$ with respect to the two multi-particle states, known as the form factor of IFT [43, 21, 20, 22, 23]. The two functions of $\rho_{L,\{\mathscr{A^{\prime}}\}_{N}}$ and $\rho_{L,\{\mathscr{A}\}_{M}}$ are Jacobian from the particle density of states of the two particle sets in the finite size system, where the detailed expressions can be found in [43]. For the diagonal matrix element where both $M=N$, and the two sets of rapidities are exactly the same, the result is given by
$\frac{\gamma_{NN}}{L}$ $\approx\sum_{i}f_{i}(t)\bigg[{}_{L}\langle\mathcal{O}_{i}\rangle_{L}$ (3)
$+\sum_{\forall\mathscr{S}\subseteq\{\mathscr{A}\}_{N}}\frac{\mathcal{F}_{\mathscr{S}\mathscr{S}}^{\mathcal{O}_{i}}(\{\vartheta+i\pi\}_{\mathscr{S}},\{\vartheta\}_{\mathscr{S}})\rho_{L,\{\mathscr{A}\}_{N}\setminus\mathscr{S}}}{\rho_{L,\{\mathscr{A}\}_{N}}}\bigg],$
with the Lüscher’s remainder as well. Here $_{L}\langle\mathcal{O}_{i}\rangle_{L}$ corresponds to the vacuum expectation value of $\mathcal{O}_{i}$, $\mathscr{S}$ is non-empty subset of $\{\mathscr{A}\}_{N}$. $\{\mathscr{A}\}_{N}\setminus\mathscr{S}$ denotes particle set excluding $\mathscr{S}$ from $\{\mathscr{A}\}_{N}$. The full information of the BCM implies that for such quantum many-body systems, we can determine all the geometric quantities associated with the BCM. Notably, the Jacobians in the denominator provide suppression behavior of order $L^{\mathcal{Z}}$ with total number of particle $\mathcal{Z}=M+N$. The algebraic decaying behaviour of the Jacobians in the limit of $L\rightarrow\infty$ [43] consequently leads to the asymptotic behaviour of the BCM elements. Thus we prove the following theorem.
Theorem: Consider time-dependent many-body quantum systems with emergent quantum integrable field theory instantaneously, if the effective adiabatic Hamiltonian follows Eq. (1), then the corresponding BCM entries with total quasi-particle numbers of $\mathcal{Z}$ follow the asymptotic algebraic behaviour $\sim L^{1-\mathcal{Z}/2}$ as the system size $L\rightarrow\infty$.
This theorem suggests that in the thermodynamic limit ($L\rightarrow\infty$) the corresponding BCM entries can be analytically determined as excitations up to two particles. This can also be understood from the quasi-particle nature of IFTs: While integrating the field operator over the entire system introduces a volume dependence, the Jacobian factor arising from the density of particle states ultimately suppresses the asymptotic behaviour of the BCM elements for multi-particle states. In the following, the theorem is applied to a time-dependent quantum Ising model, where the QGP constructed from the BCM entries is shown to play a crucial role in the MBLZT in the system.
A time-dependent quantum Ising chain with instantaneous quantum $E_{8}$ integrability.— We start from a ferromagnetic Ising chain subjected to both a periodically rotating transverse field and a static longitudinal field  [41, 38],
$\mathcal{H}(t)=-\sum_{i=1}^{L}\left(\sigma^{z}_{i}\sigma^{z}_{i+1}+\cos\omega_{0}t\sigma^{x}_{i}-\sin\omega_{0}t\sigma^{y}_{i}+h_{z}\sigma^{z}_{i}\right),$ (4)
with $\hbar=c=1$, Pauli matrices $\sigma_{i}^{\alpha}(\alpha=x,~y,~z)$ at site $i$, the length of the chain $L$, and the neighboring spin coupling $J=1$. Additionally, we require $\omega_{0}$, $h_{z}\ll 1$. It is not hard to observe that the Hamiltonian instantaneously appears as a quantum critical transverse-field Ising chain perturbed by a small longitudinal field, whose low-energy physics in the scaling limit is described by the quantum $E_{8}$ IFT,
$\mathcal{H}_{E_{8}}=\mathcal{H}_{c=1/2}-h\int\sigma(x)dx,$ (5)
where $\mathcal{H}_{c=1/2}$ is the central charge $1/2$ conformal field theory, $\sigma(x)$ is known as the spin density operator, and the coupling constant $h$ gives an exact relation with the mass of the lightest particle as $m_{1}=C_{m}h^{8/15}$, with $C_{m}=4.4049$…, and masses of all quantum $E_{8}$ particles in units of $m_{1}$ satisfy the eigenvalues of the Cartan matrix for $E_{8}$ Lie algebra [53, 22, 38, 54, 55]. In the following we set the initial state $|\psi(0)\rangle$ as the ground state of Eq. (4) at $t=0$, which in the scaling limit is the vacuum state $|0\rangle_{h}$ of the $\mathcal{H}_{E_{8}}$.
The general time-evolved wavefunction of the Hamiltonian Eq. (4) follows [41, 38] $|\psi(t)\rangle=\mathcal{U}(t)^{-1}e^{-i\mathcal{H}_{\text{eff}}t}|\psi(0)\rangle$, where the unitary transformation is given by $\mathcal{U}(t)=\exp\left\{-\frac{i}{2}\omega_{0}t\sum_{i}\sigma_{i}^{z}\right\}$, satisfying $\mathcal{U}(t)\mathcal{H}(t)\mathcal{U}(t)^{-1}=\mathcal{H}(t=0)$, and the effective Hamiltonian $\mathcal{H}_{\text{eff}}=\mathcal{U}\mathcal{H}(t)\mathcal{U}^{-1}+i\dot{\mathcal{U}}\mathcal{U}^{-1}$ follows,
$\mathcal{H}_{\text{eff}}=-\sum_{i}\left[\sigma_{i}^{z}\sigma_{i+1}^{z}+\sigma_{i}^{x}+(h_{z}-\frac{\omega_{0}}{2})\sigma_{i}^{z}\right],$ (6)
where for $h_{z}\neq{\omega_{0}}/{2}$, the low-energy physics is described by another $\mathcal{H}_{E_{8}}$ with modified longitudinal coupling. Considering the instantaneous Schrödinger equation $\mathcal{H}(t)|\psi_{n}(t)\rangle_{ins}=E_{n}(t)|\psi_{n}(t)\rangle_{ins}$, where $|\psi_{n}(t)\rangle_{ins}$ is the instantaneous eigenstate associated with the $n^{th}$ instantaneous eigenenergy $E_{n}(t)$, it is straightforward to prove [43] $E_{n}(t)=E_{n}(0)$ and $|\psi_{n}(t)\rangle_{ins}=e^{i\varphi_{n}(t)}\mathcal{U}(t)^{-1}|\psi_{n}(0)\rangle$, which implies that the instantaneous energy levels remain invariant for the Hamiltonian Eq. (4), and the instantaneous eigenstates are given by a global rotation according to $\mathcal{U}(t)$ up to an additional time-dependent overall U(1) gauge $\varphi_{n}(t)$.
Now let’s focus on the QGP $Q_{mn}$, taking the scaling limit of the model Eq. (4), the QGP follows
$Q_{nm}=\sum_{j=1}^{M}\frac{\mathcal{F}^{\mathcal{O}}_{jj}(i\pi,0)}{m_{j}\cosh\vartheta_{j}}-\sum_{i=1}^{N}\frac{\mathcal{F}^{\mathcal{O}}_{ii}(i\pi,0)}{m_{i}\cosh\vartheta_{i}}.$ (7)
As a concrete example, we consider the QGP between the vacuum state and the single $E_{8}$ particle state with zero momentum, which is given by $\mathcal{A}_{00}-\mathcal{A}_{kk}=\langle\sigma\rangle\mathcal{F}^{\sigma}_{kk}(i\pi,0)/m_{k}$ where $\langle\sigma\rangle=C_{\sigma}h^{1/15}$ with $C_{\sigma}=1.27758\cdots$ [22, 38]. The results agree with those from the boosted TCSA algorithm [47], as shown in Fig. 2 (a) for $k=1,2,3$.
The EEG in time-dependent IFT.— In the following we consider the EEG between two instantaneous eigenstates and try to find the zeros in $\Delta_{mn}(t)=0$ for any $|\psi_{m}\rangle_{ins}$ and $|\psi_{n}\rangle_{ins}$. At the zeros, the quantum adiabatic condition is significantly violated [25, 42, 26, 5], and the MBLZT process shall occur. Moreover, when $h_{z}$ decreases the EEG will cross zero and become negative, which suggests an effective swap of the high- and low- energy states instantaneously. Under this scenario, the MBLZT will be strongly enhanced because more and more higher-energy states effectively become low-energy due to the QGP. In the following we analytically solve the EEG for the ”low-energy” sector of the Hamiltonian Eq. (4), whose results are associated with the $\mathcal{H}_{E_{8}}$ [Eq. (5)].
In the scaling limit, for the low-energy sector the EEG during the time evolution can be solved analytically. Considering any two instantaneous eigenstates $|\psi_{n}(t)\rangle_{ins}$ and $|\psi_{m}(t)\rangle_{ins}$ in the model Eq. (4), the instantaneous energy gap reads $E_{mn}=\sum_{\beta=1}^{M}m_{\beta}\cosh\theta_{\beta}-\sum_{\alpha=1}^{N}m_{\alpha}\cosh\theta^{\prime}_{\alpha}$, where we have assumed that there are totally $M$ and $N$ quantum $E_{8}$ particles in the two states, respectively. With the mass ratios for the $E_{8}$ particles $\xi_{i}$ in units of $m_{1}$, the masses of the quantum $E_{8}$ particles follows $m_{\beta(\alpha)}=\xi_{\beta(\alpha)}m_{1}=\xi_{\beta(\alpha)}C_{m}h^{8/15}_{z}$, and $\theta^{(^{\prime})}_{\beta(\alpha)}$ is the corresponding rapidity. The BCM entry is given by $\gamma_{km}(t)=-\dot{\varphi}_{m}(t)\delta_{km}-\omega_{0}/2\int dx\langle\psi_{k}|\sigma(x)|\psi_{m}\rangle$. Applying the quantum $E_{8}$ form factor theory, the QGP can be organized in a more compact form as [43]
$Q_{nm}(t)=-\frac{8}{15\kappa}\left(\sum_{\beta=1}^{M}\frac{m_{\beta}}{\cosh\theta_{\beta}}-\sum_{\alpha=1}^{N}\frac{m_{\alpha}}{\cosh\theta^{\prime}_{\alpha}}\right),$ (8)
with $\kappa\equiv h_{z}/(\omega_{0}/2)$ as introduced in the introduction. Thus the EEG finally reads
$\Delta_{mn}(t)=$ $\sum_{\beta=1}^{M}m_{\beta}\cosh\theta_{\beta}-\sum_{\alpha=1}^{N}m_{\alpha}\cosh\theta^{\prime}_{\alpha}$ (9)
$-\frac{8}{15\kappa}\left(\sum_{\beta=1}^{M}\frac{m_{\beta}}{\cosh\theta_{\beta}}-\sum_{\alpha=1}^{N}\frac{m_{\alpha}}{\cosh\theta^{\prime}_{\alpha}}\right).$
Interestingly, if taking one of the states being the ground state (vacuum) of the initial states, and the other one being the single particle state with zero momentum, the zeros of EEG are given by $\kappa=8/15$, which does not depend on the type of the particles [cf. Fig. 2 (b)]. It is also the upper bound for the zeros of EEG between the ground state and the higher excited states [43]. In the quantum $E_{8}$ case, the higher excited states are readily given by single particle states with large rapidity (momentum) or multi-particle states. We show that for two-particle excitations with zero total momentum constraint, as one of the particles’ rapidity goes to infinity, the corresponding zeros of the EEG will approach 0. It is also easy to verify from Eq. (9) that the zeros for two same particle combinations with zero total momentum will collapse to the same curve as $A_{1}A_{1}$. The results are presented in Fig. 2 (c). However, if considering the zeros of EEG for any two instantaneous eigenstates, the region for zeros of EEG is beyond $\kappa=8/15$. For instance, Fig. 2 (d) shows contribution of four-particle excitations to the EEG whose zeros appear to approach $\kappa=1$.
[…]
For the time evolution governed by Eq. (4), the LE is expressed as $\mathcal{L}_{n}(t)=\langle\psi_{n}(0)|e^{-i\mathcal{H}_{\text{eff}}t}|\psi(0)\rangle$. Here, the index $n$ is used to label the energy level of the corresponding excited states. When $n$ corresponds to the ground state, we focus on the spectrum of the LE obtained through its Fourier transformation, which is given by $\mathcal{L}_{0}(\omega)=\sum_{m}|\langle\phi_{m}|\psi_{0}\rangle|^{2}\delta(\omega-E_{m})$. Here, $|\phi_{m}\rangle$ represents the eigenbasis of the $\mathcal{H}_{\text{eff}}$, and $E_{m}$ is its corresponding eigenenergy. Detailed results for $L=80$ with different $h_{z}$ from $\omega_{0}/10$ to $\omega_{0}$ are exhibited in [43]. To obtain a better understanding of the MBLZT, we introduce spectral entropy for the LE,
$S_{\text{LE}}=-\sum_{m}p_{m}\ln p_{m},$ (10)
where $p_{m}=|\langle\phi_{m}|\psi_{0}\rangle|^{2}$, and it naturally satisfies $\sum_{m}p_{m}=1$. In the context of our study, the $S_{\text{LE}}$ serves as a measure of the spectral density across the entire spectrum.
We use the TLFFA algorithm to calculate the quantities introduced above. The details of the TLFFA algorithm can be found in [43]. Specifically, in this algorithm, we truncate the free fermion basis to an energy of 10JJ above the vacuum. For different system sizes, we ensure convergent results by increasing the number of states. For instance, for $L=100$ the number truncation of states reaches $16384$ for convergence, and this also reaches our computational limit.
The results are summarized in Fig. 3. When $\kappa\gg 1$, $S_{\text{LE}}$ is small, indicating weak MBLZT. As $\kappa$ decreases, $S_{\text{LE}}$ increases continuously, suggesting an enhancement of MBLZT. When $\kappa=1$, a peak appears and increases linearly in $L$ [Fig. 3 inset (a)], which possibly implies MBLZT avalanche in thermodynamic limit. As $\kappa$ decreases further below $8/15$ upper bound of zeros of the EEG between the ground state and excited states, it is expected that the MBLZT is further strengthened, as manifested in the continuously increasing $S_{\text{LE}}$. Additional peaks appear when $\kappa<1$, signaling further strengthened MBLZT at specific $\kappa$’s. For $\kappa\rightarrow 0^{+}$, it is noticed that the increase of $S_{\text{LE}}$ becomes weaker for $L=100$ comparing to smaller $L$ [Fig. 3 (b)], which possibly is due to the truncated energy cut at 10JJ in our numerical calculation. A higher cut should provide better results, yet it exceeds our computational limit. Interestingly with $\kappa$ approaching to 1, $S_{\text{LE}}$ exhibits a hyperscaling behavior in terms of the scaling parameter $(\kappa-1)^{8/15}JL$ [Fig. 3 inset (b), $J=1$], implying emergence of universal physics near $\kappa=1$.
The entire process can also be qualitatively understood in terms of the $\mathcal{H}_{\text{eff}}$ [Eq. (6)]. Precisely at $\kappa=1$, the $\mathcal{H}_{\text{eff}}$ represents a quantum critical transverse field Ising chain with divergent density of states in the low energy sector, leading to a significant increase of the spectral entropy. When $\kappa<1$, the spin alignment in the ground state of $\mathcal{H}_{\text{eff}}$ is opposite to that of the initial ground state. This means that the initial state becomes a high-energy state with respect to the $\mathcal{H}_{\text{eff}}$, resulting in progressively stronger MBLZT as $\kappa$ decreases.
Conclusions.— In this article we lay down a theorem revealing the quantum geometry in time-dependent quantum many-body systems whose low-energy can be organized by quantum integrable field theory instantaneously. The theorem states that the BCM entries for local field densities in the system are governed by at most two-particle excitations. The result provides a geometric foundation for diabatic time evolutions mediated by the vanishing EEG due to the QGP contribution.
We then analyze a time-dependent quantum Ising chain, showing that the suppression of the EEG with decreasing longitudinal field enhances the MBLZT. Near the critical ratio $\kappa=1$, the hyperscaling behavior of the LE spectral entropy signals a sharp transition in the MBLZT. For $\kappa<1$, the numerical results show persistent entropy growth mixing with a cascade of peaks, reflecting continuous enhancement of MBLZT. Our results indicate that the LE spectral entropy can serve as a powerful diagnostic tool for diabatic time evolution and its associated MBLZT. Yet, a full physical understanding on the analytical structure of the LE spectral entropy, such as its hyperscaling behavior, linear in $L$ behavior at $\kappa=1$, and additional peaks when $\kappa<1$, requires more analytical and numerical efforts, a challenge that remains open but worth to explore further. Meanwhile it is also worth to generalize the theorem in quantum integrable lattice systems, and extend to the BCM entries for quasi-local field densities.
The time-dependent Hamiltonian in Eq. (4) is experimentally accessible in platforms such as Rydberg atomic arrays, [56, 57, 38] a promising setup for direct experimental verification of our theoretical predictions.
Figure 3: $S_{LE}$ $vs.$ $\kappa$ for different system sizes $L$’s. The $S_{LE}$ at $\kappa=1$ linearly increases with $L$ [inset (a)], along with more peaks emerging for $\kappa\lesssim 8/15$. When $\kappa$ approaches 1, the $S_{LE}$ collapses on the same curve for different $L$ [inset (b)], implying hyperscaling behavior of $S_{\text{LE}}$ w.r.t scaling parameter $(\kappa-1)^{8/15}L$.
Acknowledgements.— We thank Zongping Gong, Qicheng Tang, Thors Hans Hansson and Gabor Takács for helpful discussions. The TCSA calculation is supported by the TCSA package in [47]. The work is sponsored by National Natural Science Foundation of China Nos. 12274288, 12450004, the Innovation Program for Quantum Science and Technology Grant No. 2021ZD0301900, and the Tsung-Dao Lee Scholarship.
Data availability.—The data are available from the corresponding author upon reasonable request.